Required: Use the information in the DATA field above using cell referencing to answer the following requirements.
1. Prepare a contribution income statement for the year ignoring income tax implication.
2. Calculate the breakeven point in units.  
3. Calculate the breakeven point in sales dollars.
4. Calculate the margin of safety in sales dollars.  
5. Calculate the company's operating leverage.
6. Calculate the percentage change in profits if sales are projected to increase by: 7%
7. What if the direct labor cost per unit increases from $8 a unit to $10, what will be the new breakeven in units? Explain why it changed.
You should only have to change the direct labor in the data area and all your answers should update. Please put the direct labor cost back to the original number once you have answered the question.
8. What if the manufacturing overhead cost decreases from $90,000 to $80,000, what will be the new breakeven in units? Explain why it changed.
You should only have to change the fixed MOH in the data area and all your answers should update. Please put the fixed MOH cost back to the original number once you have answered the question.

9. The sales manager believes the company could increase sales by 700 units if advertising expenditures were increased by $20,000.  

Sales price per unit $57
Manufacturing costs: Fixed Costs Variable Costs
Direct materials $10
Direct labor 8
Variable Manufacturing overhead 5
Fixed Manufacturing overhead $90,000
Period expenses:
Variable Selling and administrative expenses 2
Fixed Selling and Administrative expenses 25,200
Totals $115,200 $25
Units sold 5,200 units
Answer #1

1)

Sales    296,400.00
Variable cost    130,000.00
Contribution Margin    166,400.00
Fixed Cost    115,200.00
Net income      51,200.00

2)Break even point in units = 115,200/(57 - 25)

=3,600 Units

3)Break even point in dollars = 115,200/0.561404

205,200

4)Margin of safety in dollars = 296,400 - 205,200 = 91,200

5) Companys Operating leverage = 166,400/51,200 = 3.25

6. Change in Profits = 3.25 x 7% = 22.75%

7.Break even point in units = 115,200/(57 - 27)

=3,840 Units

The Break even points in units has increased as contribution margin has gone down , so now company has to sell more units to recover the fixed cost

8.Break even point in units = 105,200/(57 - 25)

=3,287.50

It has gone down as the burden of fixed cost has reduced

9.

Sales    336,300.00
Variable cost    147,500.00
Contribution Margin    188,800.00
Fixed Cost    135,200.00
Net income      53,600.00